# BlockFUL: Enabling Unlearning in Blockchained Federated Learning

### v2

@@ VI-A Evaluation Metrics
AlexNet 1 $A_{D_{r}}\uparrow$ 99.99% 99.99% 99.81% 99.99% 99.99% 98.80% 90.38% 84.36% 72.50% 43.57 84.85 126.45 27.72 27.98 28.15
$A_{D_{f}}\downarrow$ 99.99% 99.99% 99.66% 0.00% 0.00% 0.00% 4.00% 3.99% 0.00%
2 $A_{D_{r}}\uparrow$ 99.99% 99.99% 99.80% 99.99% 99.99% 99.51% 81.71% 62.08% 74.26% 42.25 82.33 122.15 27.71 28.07 28.33
$A_{D_{f}}\downarrow$ 99.99% 99.99% 99.49% 0.00% 0.00% 0.00% 0.00% 0.00% 5.00%
4 $A_{D_{r}}\uparrow$ 99.99% 99.99% 99.96% 99.99% 99.99% 99.16% 72.14% 66.52% 61.03% 40.34 77.65 115.51 27.71 28.51 28.91
$A_{D_{f}}\downarrow$ 99.99% 99.99% 99.89% 0.00% 0.00% 0.00% 7.00% 0.00% 0.00%
7 $A_{D_{r}}\uparrow$ 99.99% 99.99% 99.72% 99.71% 92.72% 90.84% 90.16% 92.82% 84.03% 35.09 69.72 103.77 28.4 29.75 30.89
$A_{D_{f}}\downarrow$ 99.99% 99.99% 98.82% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
ResNet18 1 $A_{D_{r}}\uparrow$ 99.90% 99.73% 94.10% 95.36% 83.35% 94.11% 82.35% 86.08% 84.69% 261.35 515.62 763.3 27.87 28.39 28.93
$A_{D_{f}}\downarrow$ 99.99% 99.99% 94.49% 0.00% 0.00% 0.00% 0.00% 3.85% 6.27%
2 $A_{D_{r}}\uparrow$ 92.74% 92.74% 99.64% 95.67% 97.44% 91.37% 29.30% 64.45% 41.28% 236.34 465.44 697.74 28.03 28.77 29.42
$A_{D_{f}}\downarrow$ 95.75% 95.70% 98.35% 0.00% 0.00% 0.00% 5.00% 5.22% 1.81%
4 $A_{D_{r}}\uparrow$ 95.77% 98.97% 98.48% 99.96% 98.43% 95.02% 45.46% 69.25% 51.18% 175.26 353.09 532.84 28.27 29.49 30.7
$A_{D_{f}}\downarrow$ 62.73% 95.02% 99.71% 0.00% 0.00% 0.00% 9.75% 2.63% 0.12%
7 $A_{D_{r}}\uparrow$ 91.48% 87.32% 85.49% 92.59% 89.85% 98.61% 50.06% 32.05% 32.05% 115.51 223.26 330.35 28.56 30.67 32.72
$A_{D_{f}}\downarrow$ 74.46% 89.32% 91.84% 0.00% 0.00% 0.00% 5.57% 0.00% 0.00%
VI-B Datasets
The datasets utilized in the experiments are as follows:
@@ VI-B Datasets
CIFAR-10. This is a widely used dataset for image classification tasks in computer vision. It consists of 60,000 32x32 color images in 10 different classes, with 6,000 images per class. The dataset is split into 50,000 training and 10,000 testing images.
@@
Fashion-MNIST. This is a commonly used dataset for image classification tasks, similar to MNIST but more complex. The dataset comprises ten fashion categories, each comprising 60,000 grayscale images with a dimension of 28x28 pixels. The training set contains 55,000 images, while the test set includes 10,000 images.
VI-C The implementations of paradigms
@@
In our study, each traversal triggers an “unlearning” at every node, regardless of whether it has been unlearned previously or not. We employ the following unlearning methods:
Re-training. We re-train the unlearning model with a dataset containing unlearned data and the derivative models influenced by the data. We remove unlearned data from this dataset and sequentially train the model using the remaining data according to the inheritance relation.
Gradient ascent. This involves the inheritance relation where models are uploaded by the users with unlearned data. For the initial unlearning models, we calculate the difference in the model gradient before and after unlearning. Subsequently, the inheritance models utilize this gradient difference for gradient ascent. By assessing these gradients, we gain insights into the impact of re-training on the downstream models, helping us understand the unlearning process.
@@ VI-D Experimental Settings
We conduct experiments using an NVIDIA GeForce RTX 3070 GPU. We employ an untrained ResNet18 model, MobileNetV2 model, and AlexNet model, all fine-tuned on specific datasets. The to-be-unlearned classes during the process are represented as #$C_{f}$.
@@
Model Inheritance. A group of 14 users is evaluated, as depicted in Fig. 4. $user_{5}$ and $user_{7}$ have the same unlearning labels. The models highlighted in yellow are those that demonstrate accuracy in unlearned data.
@@
Blockchain. In the live chain, we use Byzantine Fault Tolerance (BFT)-based consensus to ensure data consistency and security. Additionally, we set the block packaging time to 15 seconds. This interval strikes a balance between transaction processing speed and network stability. Each block is sized at 1MB to ensure efficient network transmission and storage.
@@
Hyperparameters. In both CIFAR-10 and Fashion-MNIST setups, each model undergoes rigorous training for 100 epochs with a learning rate of 0.005.
@@
Label categories. In both CIFAR-10 and Fashion-MNIST, we apply the unlearning process to both single category labels and multiple-category labels (including combinations of 2, 4, and 7 category labels). Here, $r_{5}$, $r_{6}$, and $r_{8}$ represent the models randomly targeted for the unlearning process.
VI-E Evaluating Unlearning Performance
We evaluate the performance of re-training and gradient ascent by comparing their results for $AD_{f}$ and $AD_{r}$, which align with design goals G1 and G2, respectively. We also assess unlearning performance across single and multiple class scenarios to meet design goal G3.
We present the results in Table III for the CIFAR-10 dataset and in Table IV for the Fashion-MNIST dataset. With the Fashion-MNIST dataset, we use AlexNet for prediction. It can be observed that AlexNet’s re-training time is much shorter than that of ResNet18. This suggests that in the case of models with fewer parameters, the re-training method can achieve better unlearning effects. Moreover, the unlearning effect of AlexNet is significantly better than ResNet18 in terms of $AD_{r}$. The effectiveness of the gradient ascent method varies for different datasets and models.
VI-E1 Unlearning effectiveness of single-label unlearning
On CIFAR-10, experimental results show that $AD_{f}$ for re-training achieves 0, with an average $AD_{r}$ reaching up to 92.77%. For gradient ascent, $AD_{f}$ is around 7%, while the average $AD_{r}$ stands at 51.09%. On Fashion-MNIST, the results demonstrate that $AD_{f}$ for re-training also reaches 0, accompanied by an average $AD_{r}$ of 95.27%. For gradient ascent, $AD_{f}$ is approximately 3%, and the average $AD_{r}$ is 83.4%.
In single-label unlearning, the gradient ascent method exhibits equally impressive performance when dealing with $AD_{f}$, effectively reducing or eliminating the memory of the specified label compared to re-training the model. However, its performance on $AD_{r}$ is slightly worse, suggesting that the model may have slightly insufficient retention of other label knowledge. This indicates that while gradient ascent excels in unlearning specific labels, there is still room for improvement in maintaining the overall generalization ability of the model.
VI-E2 Unlearning effectiveness of multi-label unlearning

### v3

@@ VII-A Evaluation Metrics
AlexNet
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 90.38 84.36 72.50 43.57 84.85 126.45 27.72 27.98 28.15
$A_{D_{f}}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 4.00 3.99 0.00
2 $A_{D_{r}}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 81.71 62.08 74.26 42.25 82.33 122.15 27.71 28.07 28.33
$A_{D_{f}}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 5.00
4 $A_{D_{r}}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 72.14 66.52 61.03 40.34 77.65 115.51 27.71 28.51 28.91
$A_{D_{f}}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 7.00 0.00 0.00
ResNet18
1 $A_{D_{r}}\uparrow$ 99.90 99.73 94.10 95.36 83.35 94.11 82.35 86.08 84.69 261.35 515.62 763.3 27.87 28.39 28.93
$A_{D_{f}}\downarrow$ 99.99 99.99 94.49 0.00 0.00 0.00 0.00 3.85 6.27
2 $A_{D_{r}}\uparrow$ 92.74 92.74 99.64 95.67 97.44 91.37 29.30 64.45 41.28 236.34 465.44 697.74 28.03 28.77 29.42
$A_{D_{f}}\downarrow$ 95.75 95.70 98.35 0.00 0.00 0.00 5.00 5.22 1.81
4 $A_{D_{r}}\uparrow$ 95.77 98.97 98.48 99.96 98.43 95.02 45.46 69.25 51.18 175.26 353.09 532.84 28.27 29.49 30.7
$A_{D_{f}}\downarrow$ 62.73 95.02 99.71 0.00 0.00 0.00 9.75 2.63 0.12
TABLE III: Non-IID and IID Unlearning Performance on CIFAR-100 (ResNet18)
Methods $C_{f}$ Metrics Original (%) Re-training (%) Gradient Ascent (%) Cumulative Unlearning Time (s)
Re-training Gradient Ascent
$r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$
Label-Non-IID
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 88.40 79.34 53.86 26.45 57.41 86.42 1.27 1.79 2.37
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 11.11 5.37 0.00
2 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.44 86.16 90.32 25.14 55.36 79.14 1.47 2.14 2.38
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 4.60 14.94 4.59
3 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 90.70 62.44 67.36 23.78 51.21 73.87 1.58 2.34 2.58
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 3.41 11.32 1.50
Sample-Non-IID
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 92.81 79.55 87.06 31.64 62.51 92.64 1.84 2.35 2.94
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 1.38 0.00 0.00
2 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 81.02 79.76 91.04 29.95 59.14 88.18 1.93 2.45 3.01
$A_{D_{f}}\downarrow$ 99.98 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
3 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.80 90.09 88.89 27.67 54.61 84.48 2.05 2.55 3.19
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 2.99 2.13 1.50
IID
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 88.07 85.39 73.46 32.37 61.31 93.54 1.94 2.29 3.06
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 5.06 5.06 3.79
2 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 93.99 97.01 86.37 30.14 59.46 89.31 2.06 2.55 3.18
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
3 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 81.69 87.66 73.17 27.95 53.24 85.57 2.17 2.69 3.26
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.41 6.58 0.41
TABLE IV: Non-IID and IID Unlearning Performance on Yahoo! Answers (Bert-Base-Cased)
Methods $C_{f}$ Metrics Original (%) Re-training (%) Gradient Ascent (%) Cumulative Unlearning Time (s)
Re-training Gradient Ascent
$r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$
Label-Non-IID
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 90.41 67.42 65.63 89.63 167.54 289.56 1.78 2.98 3.04
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 1.04 7.76 7.27
2 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 87.34 62.57 56.94 70.41 156.36 220.14 2.06 3.14 3.38
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 17.35 7.99 7.01
3 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 72.04 71.75 64.68 62.78 130.21 189.87 2.58 3.34 3.58
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 7.81 16.78 14.93
Sample-Non-IID
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 85.91 71.69 61.14 107.41 214.51 321.64 1.84 2.35 2.94
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.10 0.00 0.42
2 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 95.05 89.23 90.13 94.95 185.14 303.18 1.93 3.45 3.41
$A_{D_{f}}\downarrow$ 99.98 99.99 99.99 0.00 0.00 0.00 7.47 9.83 14.38
3 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.77 77.12 77.50 86.67 161.77 285.48 2.57 3.55 3.79
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 10.63 1.84 9.92
IID
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 46.48 39.15 41.11 106.37 211.31 326.54 1.54 2.05 3.35
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 11.65 6.33 5.82
2 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 62.31 41.77 47.65 93.14 178.46 311.31 2.06 2.55 3.18
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 13.86 10.85 7.16
3 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 84.53 42.67 47.52 84.95 159.24 284.57 2.17 2.69 3.26
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 25.96 10.44 9.80
TABLE V: Non-IID Unlearning Performance on CIFAR-100 (ResNet18)
Methods $C_{f}$ Metrics Original (%) Re-training (%) Gradient Ascent (%) Cumulative Unlearning Time (s)
Re-training Gradient Ascent
$r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$
Label-Non-IID-5
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.66 97.28 99.12 26.45 57.41 86.42 1.27 1.79 2.37
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Label-Non-IID-10
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 88.55 97.98 95.52 20.64 42.51 62.64 1.05 1.35 1.94
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Label-Non-IID-15
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.01 98.11 98.71 14.37 35.31 51.54 0.89 0.94 1.06
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
TABLE VI: Non-IID Unlearning Performance on Yahoo! Answers (Bert-Base-Cased)
Methods $C_{f}$ Metrics Original (%) Re-training (%) Gradient Ascent (%) Cumulative Unlearning Time (s)
Re-training Gradient Ascent
$r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$
Label-Non-IID-1
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 88.16 72.18 83.45 104.38 206.14 314.37 1.89 3.14 3.44
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 10.28 4.09 1.07
Label-Non-IID-2
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 90.41 67.42 65.63 89.63 167.54 284.56 1.78 2.98 3.04
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 1.04 7.76 7.27
Label-Non-IID-3
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 63.04 74.86 84.77 78.34 154.31 231.36 1.68 2.57 2.67
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.94 0.63 9.73
VII-B Datasets
  • $\bullet$
@@ VII-B Datasets
    CIFAR-10: Comprises 60,000 32x32 color images in 10 different classes, with 6,000 images per class. The dataset is split into 50,000 training and 10,000 testing images.
  • $\bullet$
@@
    Fashion-MNIST: Comprises 10 fashion categories, each comprising 60,000 grayscale images with a dimension of 28x28 pixels. The training set contains 55,000 images, while the test set includes 10,000 images.
  • $\bullet$
    Yahoo! Answers: Contains approximately 1.4 million entries across 10 categories, offering a valuable resource for natural language processing (NLP) tasks.
  • $\bullet$
    TinyImageNet: The dataset contains 200 classes, with 500 images per class in the training set and an additional 10,000 images in the test set.
VII-C Implementation of Unlearning Paradigms
@@
Each traversal triggers an “unlearning” at every node, regardless of whether it has been unlearned previously or not. We employ the following unlearning methods:
Re-training (Sequential). The model is re-trained sequentially on the remaining data after removing the unlearned data, following the inheritance path.
Gradient ascent (Parallel). We calculate gradient differences from initial unlearning and apply them to inherited models in parallel, enabling efficient unlearning and analysis of its downstream effects.
@@ VII-D Experimental Settings
We conduct experiments using an NVIDIA GeForce RTX 3070 GPU (Tables I and II) and an RTX 4090 GPU (Tables III-X and Section VII-F). We employ an untrained ResNet18 model, MobileNetV2 model, Bert-Base-Cased model, GCNN model, and AlexNet model, all fine-tuned on specific datasets. The to-be-unlearned classes during the process are represented as $C_{f}$.
  • $\bullet$
@@
    Model Inheritance: A group of 14 users is evaluated, as depicted in Fig. 4, where $user_{5}$ and $user_{7}$ have the same unlearning labels. The models highlighted in yellow are those that demonstrate accuracy in unlearned data.
  • $\bullet$
@@
    Blockchain: In the live chain, we use Byzantine Fault Tolerance (BFT)-based consensus to ensure data consistency and security. Additionally, we set the block packaging time to 15 seconds. This interval strikes a balance between transaction processing speed and network stability. Each block is sized at 1MB to ensure efficient network transmission and storage.
  • $\bullet$
@@
    Hyperparameters: In both CIFAR-10 and Fashion-MNIST setups, each model undergoes rigorous training for 100 epochs with a learning rate of 0.005.
  • $\bullet$
@@
    Label categories: We apply the unlearning process to both single and multiple category labels (such as combinations of 2, 4, etc., category labels). Here, $r_{5}$, $r_{6}$, and $r_{8}$ represent the models randomly targeted for the unlearning process.
VII-E Evaluating Unlearning Performance
We evaluate re-training and gradient ascent by comparing $AD_{f}$ and $AD_{r}$, corresponding to design goals G1 and G2, while assessing single- and multi-label unlearning to address G3. Tables I and II report results on CIFAR-10 and Fashion-MNIST using AlexNet, which shows shorter re-training time and better $AD_{r}$ than ResNet18, indicating that smaller models may unlearn more effectively via re-training.
VII-E1 Unlearning effectiveness of single-label unlearning
On CIFAR-10 and Fashion-MNIST, re-training consistently achieves perfect unlearning ($AD_{f}$=0) with high retention ($AD_{r}$ up to 92.77% and 95.27%, respectively). In contrast, gradient ascent yields moderate unlearning ($AD_{f}$ around 7% and 3%) but lower retention ($AD_{r}$of 51.09% and 83.4%). In single-label unlearning, gradient ascent matches re-training in $AD_{f}$, effectively forgetting the target label. However, its lower $AD_{r}$ indicates weaker retention of other labels.
VII-E2 Unlearning effectiveness of multi-label unlearning